Grade 12th passGeneral Physics

Two ice skaters of weight 60 kg and 50 kg are holding the two ends of a rope. The rope
is taut. The 60 kg man pulls the rope with 20 N force. What will be the force exerted by
the rope on the other person? What will be their respective acceleration?

Ssugan Sugan
10 Years agoGrade 12th pass

3 Answers

Riddhish Bhalodia
10 Years ago
The tension is propogated and hence the force across the string remains as 20N which is also the force applied on the second man
The respective acceleration will be
1: 20/60 = 1/3
2: 20/50 = 2/5
Riddhish Bhalodia
10 Years ago
The tension is propogated and hence the force across the string remains as 20N which is also the force applied on the second man
The respective acceleration will be
1: 20/60 = 1/3
2: 20/50 = 2/5
Shoba
7 Years ago
According to newtons 3 law
Every force has an equal and opposie force
So
 
F1=-F2
So the force is applied on second man is  -20N